### v1

# AdagradLSPIA: Integrating adaptive optimization into least squares progressive iterative approximation

Abstract
This paper introduces the Adaptive Gradient Least Squares Progressive iterative Approximation (AdagradLSPIA), an accelerated version of the Least Squares Progressive Iterative Approximation (LSPIA) method, enhanced with adaptive optimization techniques inspired by the adaptive gradient (Adagrad) algorithm. By dynamically adjusting weights using historical gradient information, AdagradLSPIA achieves faster convergence compared to the standard LSPIA method. The method’s effectiveness is demonstrated through tensor product B-spline surface fitting, where it consistently outperforms LSPIA in terms of accuracy, computational efficiency, and robustness to variations in global weight selection. Theoretical analysis confirms its convergence guarantees, and experimental results highlight its suitability for complex surface fitting tasks in geometric modeling and reverse engineering.
keywords:
Geometric iterative method , Least squares progressive iterative approximation , Adaptive optimization , Adagrad , Surface fitting
[…]
url]https://sajavicius.github.io/
1 Introduction
In geometric design, data fitting, and reverse engineering, creating accurate and efficient representations of complex shapes and surfaces from given data points is a fundamental task. This is particularly important in applications such as computer graphics, CAD/CAM systems, and digital manufacturing, where precise surface fitting is essential for quality modeling, simulation, and fabrication. Among the various methods developed for these purposes, Least Squares Progressive Iterative Approximation (LSPIA) has gained significant attention due to its simplicity and effectiveness in handling large sets of data points.
LSPIA, initially introduced by Deng and Lin [1], is a geometric iterative method (GIM) that improves on the standard Progressive Iterative Approximation (PIA) [2] by minimizing the least squares error between the data points and the approximate surface. The method constructs a sequence of fitting patches by iteratively updating the control points of the surface, thereby reducing the fitting error at each step. This approach has been widely adopted for various geometric modeling tasks, including mesh generation, solid modeling, and reverse engineering [3]. Despite its effectiveness, the standard LSPIA method relies on a constant global weight to scale adjusting vectors, which can limit the convergence speed and accuracy, especially when dealing with data sets that exhibit significant variability.
Recent research has introduced several enhancements to the LSPIA method, emphasizing improvements in convergence speed and computational efficiency. For instance, memory-augmented methods like MLSPIA [4] have been proposed, demonstrating superior performance through advanced weighting mechanisms and relaxation strategies. The asynchronous LSPIA (ALSPIA) [5] method significantly reduces iteration counts and computational time compared to the standard LSPIA, particularly in large-scale problems, by leveraging asynchronous updates to the adjusting vectors. Another notable advancement is the introduction of momentum techniques in LSPIA, as seen in the PmLSPIA and NmLSPIA methods [6], which utilize Polyak’s and Nesterov’s momentum to achieve faster convergence by better determining the search direction. High-order iterative methods of matrix inverse computation have also been employed to further optimize the convergence rates and computational efficiency of LSPIA [7, 8, 9]. These advancements underscore the continuous evolution of LSPIA, highlighting its adaptability and efficacy in addressing complex geometric modeling tasks.
One way to improve the performance of iterative methods like LSPIA is by adopting adaptive optimization techniques. Adaptive optimization methods, such as the Adaptive Gradient (Adagrad) algorithm [10], have become popular in machine learning due to their ability to dynamically adjust learning rates based on historical gradient information (see also [11, 12]). This adaptability allows for faster and more stable convergence, particularly when dealing with data where gradient magnitudes vary significantly across parameters. Inspired by this, we propose an accelerated version of the standard LSPIA, named Adaptive Gradient Least Squares Progressive Iterative Approximation (AdagradLSPIA).
The AdagradLSPIA method integrates the adaptive weighting mechanism of Adagrad with the geometric iterative framework of LSPIA. Unlike the standard approach, where a fixed global weight is used, AdagradLSPIA assigns separate adaptive weights to each adjusting vector based on accumulated gradient information from previous iterations. This not only enhances the convergence rate by allowing larger weights in smoother regions of the surface but also provides more stability when handling complex geometric features.
The remainder of this paper is organized as follows. In section 2, the formulation of the AdagradLSPIA method is presented, which includes a detailed discussion on its adaptive weighting mechanism, the underlying geometric iterative procedure, and the theoretical convergence analysis. This is followed by an extensive experimental study in section 3 that demonstrates the effectiveness of the proposed method. Finally, section 4 concludes the paper with a summary of the findings and suggestions for future research.
2 Adaptive gradient least squares progressive iterative approximation
In this section, the Adaptive Gradient Least Squares Progressive Iterative Approximation (AdagradLSPIA) method is introduced, which integrates adaptive optimization techniques into the traditional LSPIA framework. We begin by formulating the data fitting problem and geometric iterative method for its solution, followed by a detailed explanation of the adaptive weighting mechanism that drives the convergence of the method. We also discuss the key aspects of hyperparameter tuning and convergence.
2.1 Problem formulation
The goal of the AdagradLSPIA method is to fit a smooth curve or surface to a given set of data points by minimizing the least squares error.
Let $\mathcal{I}$ and $\mathcal{J}$ be appropriately chosen (multi-)index sets such that $\lvert\mathcal{I}\rvert\leq\lvert\mathcal{J}\rvert$. Consider a set of data points $\{\mathbf{q}_{j}\}_{j\in\mathcal{J}}\subset\mathbb{R}^{d}$ (where $d=2,3$) with corresponding parameter values $\{\mathbf{t}_{j}\}_{j\in\mathcal{J}}$ in a suitable parameter space $\Omega\subset\mathbb{R}$ (or $\mathbb{R}^{2}$ for surfaces). The fitting patch, i.e. a curve or surface that approximates the given data points, is represented as
$\mathbf{c}(\mathbf{t})=\sum_{i\in\mathcal{I}}{\mathbf{p}_{i}B_{i}(\mathbf{t})},\quad\mathbf{t}\in\Omega,$
where $\mathbf{p}_{i}\in\mathbb{R}^{d}$ are the control points to be determined, and $B_{i}(\mathbf{t})$ are blending functions (e.g., B-spline or NURBS basis functions).
The considered fitting problem can be formulated as the least squares problem
$E(\mathbf{p})=\frac{1}{2}\sum_{j\in\mathcal{J}}{\lVert\mathbf{q}_{j}-\mathbf{c}(\mathbf{t}_{j})\rVert_{2}^{2}}\rightarrow\min$ ((1))
where $E(\mathbf{p})$ is the objective function (least squares error), and $\mathbf{p}=[\mathbf{p}_{i}]_{i\in\mathcal{I}}^{T}$ is the vector containing the control points of the fitting patch $\mathbf{c}(\mathbf{t})$.
2.2 Geometric iterative procedure with adaptive weights
To solve the above problem, an iterative procedure is employed. This procedure iteratively adjusts control points and constructs a sequence of fitting patches $(\mathbf{c}^{(k)}(\mathbf{t}))_{k=0}^{\infty}$.
The procedure is initialized by selecting an initial set of control points $\mathbf{p}_{i}^{(0)}$ which are used to construct the initial fitting patch
$\mathbf{c}^{(0)}(\mathbf{t})=\sum_{i\in\mathcal{I}}{\mathbf{p}_{i}^{(0)}B_{i}(\mathbf{t})},\quad\mathbf{t}\in\Omega.$
Choosing the initial control points approporiately is important to ensure the rapid convergence of the iterative procedure [1].
At each iteration, the difference vectors between the data points and the current fitting patch $\mathbf{c}^{(k)}(\mathbf{t})$ are calculated as
$\boldsymbol{\delta}_{j}^{(k)}=\mathbf{q}_{j}-\mathbf{c}^{(k)}(\mathbf{t}_{j}),\quad j\in\mathcal{J}.$
These difference vectors measure the discrepancies between the data and the approximated surface. The adjusting vectors, which determine how the control points are updated, are computed as
$\boldsymbol{\varDelta}_{i}^{(k)}=-\sum_{j\in\mathcal{J}}{B_{i}(\mathbf{t}_{j})\boldsymbol{\delta}_{j}^{(k)}},\quad i\in\mathcal{I}.$
The control points are then updated as
$\mathbf{p}_{i}^{(k+1)}=\mathbf{p}_{i}^{(k)}-\mu_{i}^{(k+1)}\boldsymbol{\varDelta}_{i}^{(k)},\quad i\in\mathcal{I},$
where $\mu_{i}^{(k+1)}$ are the adaptive weights that dynamically adjust during the iterations. The sequence of fitting patches is thus generated as
$\mathbf{c}^{(k+1)}(\mathbf{t})=\sum_{i\in\mathcal{I}}{\mathbf{p}_{i}^{(k+1)}B_{i}(\mathbf{t})},\quad\mathbf{t}\in\Omega.$
The control point update rule can be expressed as
$\mathbf{p}^{(k+1)}=\mathbf{p}^{(k)}-\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)},$ ((2))
where $\mathbf{p}^{(k)}=[\mathbf{p}_{i}^{(k)}]_{i\in\mathcal{I}}^{T}$, $\mathbf{W}^{(k+1)}=\operatorname{diag}\left(\mu_{i}^{(k+1)}\right)_{i\in\mathcal{I}}$, and $\boldsymbol{\Delta}^{(k)}=[\boldsymbol{\varDelta}_{i}^{(k)}]_{i\in\mathcal{I}}^{T}$.
Rios and Jüttler [13] established a connection between LSPIA and stochastic gradient descent by demonstrating that the adjusting vectors used in LSPIA correspond to components of the gradient vector of the objective function. Considering the objective function $E(\mathbf{p})$ defined by ((1)), the adjusting vectors $\boldsymbol{\varDelta}_{i}^{(k)}$ align with the gradient components, such that $\nabla E(\mathbf{p}^{(k)})=\boldsymbol{\Delta}^{(k)}$.
2.3 Adaptive weighting mechanism
A critical enhancement in AdagradLSPIA compared to standard LSPIA is the introduction of an adaptive weighting mechanism inspired by the Adagrad optimization algorithm. Unlike standard LSPIA, which uses a fixed global weight to scale the adjusting vectors, AdagradLSPIA employs separate adaptive weights for each control point adjustment. These weights are defined as
$\mu^{(k+1)}_{i}=\frac{\mu}{\sqrt{\epsilon+v^{(k+1)}_{i}}},\quad i\in\mathcal{I},$ ((3))
where:
  • $\mu$ is a global weight (also known as the learning rate in the context of machine learning) that controls the overall weight (step size);
  • $\epsilon>0$ is a small constant that ensures numerical stability, preventing division by zero;
  • $v^{(k+1)}_{i}$ is the accumulated gradient information, calculated as
    $v^{(0)}_{i}=0,\quad v^{(k+1)}_{i}=v^{(k)}_{i}+\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2},\quad i\in\mathcal{I}.$
The adaptive weights $\mu^{(k+1)}_{i}$ allow the method to take larger steps in directions where the gradients are smaller, thereby accelerating convergence in smoother regions of the fitting patch. Conversely, the weights decrease when the gradients are large, ensuring that the method remains stable and does not overshoot during the iterative process.
2.4 Hyperparameter tuning
The performance of the AdagradLSPIA method depends on the appropriate selection of hyperparameters, particularly $\mu$ and $\epsilon$. The global weight $\mu$ determines the scale of adjustments to the control points, and it is important to balance this parameter to avoid overly aggressive updates, which can lead to oscillations, or overly conservative updates, which may slow down convergence. In practice, $\mu$ can be selected through cross-validation or empirical testing.
The stability parameter $\epsilon$ is usually set to a small value (e.g. $10^{-8}$) to avoid division by zero in the calculation of $\mu^{(k+1)}_{i}$. Although $\epsilon$ does not have a significant impact on convergence behavior, it can be adjusted based on the scale of the gradients to improve numerical stability.
2.5 Convergence analysis
The convergence analysis of the AdagradLSPIA method is crucial to ensure that the algorithm consistently minimizes the least squares objective and reaches an optimal solution. Given its adaptive weights, this section presents key results showing that the method ensures a steady decrease in the error while maintaining stability. These findings confirm that AdagradLSPIA converges to the solution of the least squares problem under standard conditions of convexity and smoothness.
Recall that the objective function $E(\mathbf{p})$ is convex. Moreover, the gradient $\nabla E(\mathbf{p})$ is Lipschitz continuous, i.e., there exists a constant $L>0$ such that
$\left\lVert\nabla E(\mathbf{p}_{1})-\nabla E(\mathbf{p}_{2})\right\rVert\leq L\left\lVert\mathbf{p}_{1}-\mathbf{p}_{2}\right\rVert,\quad\forall\mathbf{p}_{1},\mathbf{p}_{2}.$
Before proving the convergence of the AdagradLSPIA method, it is essential to show that each iteration leads to a decrease in the objective function $E(\mathbf{p})$. subsection 2.5 establishes this descent property, ensuring that the adaptive weights and updates consistently reduce $E(\mathbf{p})$, which is crucial for guaranteeing overall convergence.
[…]
For a convex objective function $E(\mathbf{p})$ with a Lipschitz continuous gradient, there exists a choice of the adaptive weight matrix $\mathbf{W}^{(k+1)}$ such that
$E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})-\frac{1}{2}\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}.$
Proof.
To derive the inequality, we start with the Taylor expansion of the convex function $E(\mathbf{p})$ around the point $\mathbf{p}^{(k)}$. Using a first-order Taylor expansion, we can approximate $E(\mathbf{p}^{(k+1)})$ as
$E(\mathbf{p}^{(k+1)})\approx E(\mathbf{p}^{(k)})+\boldsymbol{\Delta}^{(k)T}(\mathbf{p}^{(k+1)}-\mathbf{p}^{(k)}).$
Since the gradient $\nabla E(\mathbf{p})$ is Lipschitz continuous, we can include a quadratic term to bound the function more accurately. Specifically, we have
$E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})+\boldsymbol{\Delta}^{(k)T}(\mathbf{p}^{(k+1)}-\mathbf{p}^{(k)})+\frac{L}{2}\lVert\mathbf{p}^{(k+1)}-\mathbf{p}^{(k)}\rVert^{2}.$
From the update rule ((2)), we get
$\mathbf{p}^{(k+1)}-\mathbf{p}^{(k)}=-\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}.$
By substituting this into the bound for $E(\mathbf{p}^{(k+1)})$, we obtain
$E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})-\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}+\frac{L}{2}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}.$
The goal is to demonstrate that the negative term, $-\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\nabla E(\mathbf{p}^{(k)})$, outweighs the positive term $\frac{L}{2}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$. This will confirm that $E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})$ and, ideally, show a net decrease.
To proceed, let $\boldsymbol{\Delta}^{(k)}=\nabla E(\mathbf{p}^{(k)})$. We can rewrite the inequality as
$E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})-\frac{1}{2}\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}-\left(\frac{1}{2}\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}-\frac{L}{2}\lvert|\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rvert|^{2}\right).$
Now, we need to control the term inside the parentheses. The matrix $\mathbf{W}^{(k+1)}$ is designed to adapt based on historical gradient information, with each diagonal element $\mu_{i}^{(k+1)}$ defined as in ((3)). Substituting this form of $\mathbf{W}^{(k+1)}$, we observe that
$\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}=\sum_{i\in\mathcal{I}}\mu_{i}^{(k+1)}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2},$
and
$\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}=\sum_{i\in\mathcal{I}}\mu_{i}^{(k+1)2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}.$
To ensure that the negative term dominates, we need
$\frac{\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}}{\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}}\geq L.$
Simplifying, we want
$\frac{\sum_{i\in\mathcal{I}}\mu_{i}^{(k+1)}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}}{\sum_{i\in\mathcal{I}}\mu_{i}^{(k+1)2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}}\geq L.$
We can effectively control the weights $\mu_{i}^{(k+1)}$ through the following considerations:
  • Adaptive scaling: Each weight $\mu_{i}^{(k+1)}$ is inversely proportional to $\sqrt{\epsilon+v_{i}^{(k+1)}}$, where $v_{i}^{(k+1)}$ is the accumulated sum of squared gradients for the $i$-th component up to iteration $k$. This design ensures that if $\boldsymbol{\varDelta}_{i}^{(k)}$ is consistently large, $v_{i}^{(k+1)}$ grows, and $\mu_{i}^{(k+1)}$ decreases. Therefore, the method automatically reduces the weight in directions where the gradient is persistently high, preventing excessive updates that could destabilize convergence.
  • Balancing weights: The choice of $\mu$ affects the overall scale of $\mu_{i}^{(k+1)}$. By setting $\mu\leq\frac{1}{L}$, we ensure that the weights $\mu_{i}^{(k+1)}$ are not too large, preventing the method from overshooting and ensuring that the term $\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ does not dominate. This careful scaling helps balance the adjustments and ensures the desired decrease in $E(\mathbf{p})$.
  • Effect of $\epsilon$: The constant $\epsilon$ prevents $\mu_{i}^{(k+1)}$ from becoming too small too quickly, especially during the early stages when $v_{i}^{(k+1)}$ may still be relatively low. This helps in taking more confident steps early on, promoting faster initial convergence. As $v_{i}^{(k+1)}$ accumulates over iterations, $\epsilon$ becomes less influential, and the adaptive mechanism takes over, ensuring stability.
Consequently, if $\mu$ is chosen such that $\mu\leq\frac{1}{L}$, then
$\frac{\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}}{\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}}\approx\frac{1}{\mu}\geq L,$
satisfying the desired inequality and proving that
$E(\mathbf{p}^{(k+1)})\leq E(\mathbf{p}^{(k)})-\frac{1}{2}\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}.$
Thus, this guarantees that the objective function decreases at each iteration, confirming the descent property and proving subsection 2.5. ∎
To ensure the stability and convergence of the AdagradLSPIA method, it is important to verify that the adaptive weights do not become excessively small or large over iterations. subsection 2.5 addresses this by proving that the accumulated gradient information keeps the weights $\mu_{i}^{(k)}$ balanced, preventing erratic behavior in the updates, and ensuring that the method can continue to make steady progress toward minimizing the objective function.
[…]
The accumulation of squared gradients $v_{i}^{(k)}$ ensures that adaptive weights $\mu_{i}^{(k)}$ do not become too small, thus preventing instability. Specifically:
$\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}<\infty.$
Proof.
Let $\boldsymbol{\varDelta}_{i}^{(k)}$ represent the $i$-th component of the gradient $\boldsymbol{\Delta}^{(k)}$. We need to show that the series $\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ converges, which implies that the sequence $\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert$ will approach zero as $k\to\infty$.
First, note that the accumulated gradient $v_{i}^{(k+1)}$ grows as more gradients are added:
$v_{i}^{(k+1)}=v_{i}^{(0)}+\sum_{t=1}^{k}\lVert\boldsymbol{\varDelta}_{i}^{(t)}\rVert^{2}.$
Since $\lVert\boldsymbol{\varDelta}_{i}^{(t)}\rVert^{2}$ is non-negative for all $t$, $v_{i}^{(k+1)}$ is a non-decreasing sequence that becomes larger as $k$ increases. This accumulation plays a key role in adjusting the adaptive weight $\mu_{i}^{(k+1)}$.
The adaptive weight $\mu_{i}^{(k+1)}$ is defined by ((3)). As $v_{i}^{(k+1)}$ accumulates, the denominator $\sqrt{\epsilon+v_{i}^{(k+1)}}$ in ((3)) increases, leading to a gradual decrease in $\mu_{i}^{(k+1)}$. This ensures that the weights for directions with consistently large gradients shrink, preventing instability and oscillations.
Now consider the sum
$\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}=\sum_{i\in\mathcal{I}}\mu_{i}^{(k+1)2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}.$
Substituting the expression for $\mu_{i}^{(k+1)}$, we have
$\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}=\sum_{i\in\mathcal{I}}\frac{\mu^{2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}}{\epsilon+v_{i}^{(k+1)}}.$
To ensure that the series $\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ converges, we analyze the behavior of each term. Since $v_{i}^{(k+1)}$ is non-decreasing, it grows over iterations, leading to larger denominators $\epsilon+v_{i}^{(k+1)}$. This causes the term $\frac{\mu^{2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}}{\epsilon+v_{i}^{(k+1)}}$ to decrease, especially for parameters where $\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}$ remains consistently large over iterations. Consequently, adaptive weights effectively reduce the contribution of gradients that accumulate over time, ensuring stability.
Given that $v_{i}^{(k+1)}$ grows without bound for parameters that experience persistent gradients, the denominator $\epsilon+v_{i}^{(k+1)}$ will eventually dominate, leading to
$\lim_{k\to\infty}\frac{\mu^{2}\lVert\boldsymbol{\varDelta}_{i}^{(k)}\rVert^{2}}{\epsilon+v_{i}^{(k+1)}}=0.$
This ensures that each term $\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ decreases over time, contributing less to the overall sum as $k$ increases.
Therefore, the series $\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ is convergent, implying that $\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert\rightarrow 0$ as $k\rightarrow\infty$. This convergence guarantees that the effective weights shrink appropriately, maintaining stability, and preventing erratic behavior. Hence, the adaptive weights $\mu_{i}^{(k+1)}$ do not become excessively small or large, allowing the AdagradLSPIA method to make steady progress toward minimizing the objective function, as required.
This completes the proof of subsection 2.5. ∎
Before proving the overall convergence of the AdagradLSPIA method, it is important to establish that each iteration not only decreases the objective function (as shown in subsection 2.5) but also maintains stability in the adaptive weights $\mu_{i}^{(k+1)}$ (as ensured by subsection 2.5). Together, these two properties allow us to demonstrate that the method steadily approaches a global minimum. The following theorem combines these results to confirm the convergence behavior and rate of the AdagradLSPIA algorithm.
Theorem 1 (Convergence of AdagradLSPIA).
The AdagradLSPIA method converges to a global minimum of $E(\mathbf{p})$ at a sublinear rate $\mathcal{O}(1/\sqrt{K})$.
Proof.
Using subsection 2.5, we have
$E(\mathbf{p}^{(k)})-E(\mathbf{p}^{(k+1)})\geq\frac{1}{2}\boldsymbol{\Delta}^{(k)T}\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}.$
Summing over iterations $k=1,\ldots,K$, we get
$\sum_{k=1}^{K}\left(E(\mathbf{p}^{(k)})-E(\mathbf{p}^{(k+1)})\right)\geq\frac{1}{2}\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}.$
This can be simplified as
$E(\mathbf{p}^{(1)})-E(\mathbf{p}^{(K+1)})\geq\frac{1}{2}\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}.$
Since $E(\mathbf{p})$ is lower-bounded (e.g., $E(\mathbf{p})\geq 0$), this inequality ensures that the series $\sum_{k=1}^{K}\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert^{2}$ converges. Consequently, $\lVert\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}\rVert\rightarrow 0$ as $k\rightarrow\infty$. Because $\mathbf{W}^{(k+1)}$ is diagonal and positive-definite, this implies
$\lVert\boldsymbol{\Delta}^{(k)}\rVert\rightarrow 0\quad\text{as}\quad k\rightarrow\infty,$
which satisfies the first-order optimality condition $\nabla E(\mathbf{p})=0$.
To further establish convergence, it is essential to ensure that the weights $\mu_{i}^{(k+1)}$ within $\mathbf{W}^{(k+1)}$ do not behave erratically, which could lead to instability or excessively small weights. subsection 2.5 confirms that the accumulation of gradient information $v_{i}^{(k+1)}$ ensures that the weights $\mu_{i}^{(k+1)}$ remain bounded and balanced. This prevents the effective weights from vanishing or exploding, allowing the optimization process to proceed steadily. As $k$ increases, the adaptive weights adjust to the magnitude of the gradients, ensuring that $\mathbf{W}^{(k+1)}\boldsymbol{\Delta}^{(k)}$ diminishes smoothly over time.
Since $E(\mathbf{p})$ is convex, the point $\mathbf{p}^{*}$ where $\nabla E(\mathbf{p}^{*})=0$ is a global minimum. Thus, the AdagradLSPIA method converges to the global minimum $\mathbf{p}^{*}$. The sublinear rate $\mathcal{O}(1/\sqrt{K})$ follows from the standard convergence analysis for adaptive gradient methods, as shown in previous work (e.g. [10]). This convergence rate is supported by the mechanism described in subsection 2.5, which ensures that adaptive weights decrease appropriately to allow steady and consistent progress toward the minimum. ∎
The convergence analysis of AdagradLSPIA demonstrates that the method effectively minimizes the least squares objective function. The adaptive weights $\mu_{i}^{(k+1)}$ ensure stability and convergence, while the use of accumulated gradient information helps the method achieve a sublinear convergence rate. This analysis highlights the robustness and efficiency of AdagradLSPIA compared, for example, to the standard LSPIA method.
3 Experimental study
This section details the implementation of the AdagradLSPIA method and presents the experimental results of its comparison with the standard LSPIA method in the context of B-spline surface fitting.
3.1 Implementation details
The experiments utilize a topologically rectangular set of $251\times 251$ data points representing an automobile hood that features a combination of smooth regions, sharp transitions, and subtle design details (see Figure 1). Parameter values for the data points were assigned using a uniform parameterization method in both parametric directions.
A bicubic tensor product B-spline surface with $50\times 50$ control points was employed for fitting. The knot vectors to define the B-spline basis functions were constructed using the averaging technique (see [14, 15]). For initialization, all control points were set to zero.
Fig. 1: Data points representing an automobile hood model.
The fitting error was computed at each iteration as the sum of the squared Euclidean distances between the data points and their corresponding points on the fitting surface. The iterative process terminated either when the absolute difference between consecutive fitting errors was less than $10^{-7}$ or when the maximum of $10^{3}$ iterations was reached.
3.2 Results and discussion
Influence of global weight
The initial set of experiments investigated how global weight affected the performance of both methods. LSPIA was tested with 200 global weight values sampled uniformly from the interval $(0,0.2]$, while AdagradLSPIA was examined over $(0,20]$. The results (Figure 2) show that LSPIA converged only within a limited range of global weights; outside this range, it reached the maximum iteration limit without achieving the desired accuracy. In contrast, AdagradLSPIA converged consistently across the entire tested range, demonstrating its robustness to variations in global weight.
Optimized global weights
From the experiments, global weight values yielding minimal fitting error, minimal elapsed time, and minimal iterations were identified for both methods. Table 1 summarizes these results, showing that AdagradLSPIA achieved better accuracy, reduced computational time, and required fewer iterations compared to LSPIA. For example, the minimal fitting error for AdagradLSPIA was $8.75530\times 10^{-7}$, which is significantly lower than $1.78478\times 10^{-6}$ achieved by LSPIA.
Convergence behavior
Convergence plots (Figure 3) for the optimized global weights reveal sublinear convergence for both methods, with AdagradLSPIA consistently outperforming LSPIA in reducing the fitting error per iteration and per unit of elapsed time. AdagradLSPIA demonstrated faster error reduction, particularly in the initial stages, reflecting the benefits of its adaptive weighting mechanism.
Surface quality
The final fitting surfaces (Figure 4), obtained using the global weights corresponding to the minimal fitting errors, and their reflection lines (Figure 5) provide a qualitative comparison of the results. The reflection lines show the ability of AdagradLSPIA to capture geometric details and maintain surface smoothness.
Fig. 2: Comparison of the LSPIA and AdagradLSPIA methods across a range of global weight values. The plots depict the variation of fitting error, elapsed time (in s), and the number of iterations required to converge, highlighting the broader convergence region and superior performance of AdagradLSPIA.
Table 1: Performance comparison of the LSPIA and AdagradLSPIA methods for the automobile hood model. The table includes the initial fitting error, final fitting errors, elapsed times (in s), and the number of iterations required for global weights corresponding to minimal fitting error, elapsed time, and iterations.
Method Global weight Fitting error Elapsed time Iterations
Initial fitting surface
7.03910E+04
Minimal fitting error
LSPIA 0.017231 1.78478E-06 0.711779 272
AdagradLSPIA 2.200000 8.75530E-07 0.150992 58
Minimal elapsed time
LSPIA 0.016849 3.32523E-06 0.392019 154
AdagradLSPIA 19.500000 1.07609E-06 0.140959 56
Minimal number of iterations
LSPIA 0.016754 3.47095E-06 0.398633 153
AdagradLSPIA 16.200000 1.15355E-06 0.145413 56
Fig. 3: Convergence plots for the LSPIA and AdagradLSPIA methods using optimized global weight values. The plots show the reduction in fitting error over elapsed time and iterations, demonstrating the faster convergence of AdagradLSPIA. (For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.)
Fig. 4: The final fitting surfaces generated by the LSPIA and AdagradLSPIA methods. The surfaces were obtained using the respective global weights that minimize the fitting error.
Fig. 5: Reflection lines (zebra mappings) on the fitting surfaces generated by the LSPIA and AdagradLSPIA methods. These lines demonstrate the ability of AdagradLSPIA to capture geometric details and maintain surface smoothness.
3.3 Summary of results
The experiments conclusively demonstrate the advantages of AdagradLSPIA over the standard LSPIA method. AdagradLSPIA consistently achieved higher accuracy, faster convergence, and improved surface quality. Its robustness to global weight variations and adaptability to the geometric characteristics of the data make it a superior choice for surface fitting tasks.
4 Concluding remarks
This paper introduced the Adaptive Gradient Least Squares Progressive Iterative Approximation (AdagradLSPIA) method, a significant enhancement of the standard LSPIA technique for surface fitting. By integrating the Adagrad optimization algorithm into the geometric iterative framework, AdagradLSPIA dynamically adjusts weights based on accumulated gradient information, resulting in more stable and efficient convergence.
Theoretical analysis showed that AdagradLSPIA maintains the guaranteed convergence of LSPIA while achieving a sublinear convergence rate of $\mathcal{O}(1/\sqrt{K})$. The experimental results confirmed the practical advantages of the method, including faster convergence, reduced computational cost, and improved accuracy compared to the standard approach. In particular, AdagradLSPIA exhibited superior performance in a range of global weight values and consistently produced smooth and accurate fitting surfaces, as validated by reflection line analysis.
In addition to its promising results, AdagradLSPIA opens avenues for further research. The integration of other advanced adaptive optimizers, such as Adam [16], Adadelta [17], Nadam [18], hypergradient descent [19], or AMSGrad [20], could potentially improve the efficiency and robustness of the method. Future work may also extend the method to handle more complex geometries. Overall, AdagradLSPIA represents a robust and efficient tool for surface fitting, with significant potential for applications in geometric modeling, CAD, and reverse engineering.